OpenStudy (anonymous):

anybody prove this 0=2?

OpenStudy (anonymous):

2 doesnt equal zero

Parth (parthkohli):

2 equals 0 in a universe where people don't know Mathematics.

OpenStudy (anonymous):

@ParthKohli Funny one.......

OpenStudy (anonymous):

we cant divide by zero

OpenStudy (anonymous):



OpenStudy (anonymous):

lol ... we cant divide by zero

OpenStudy (anonymous):

hey @SOMU.kvp don't you hae mathematics in your universe....... Its not possible.....

Parth (parthkohli):

I'm calling the police for you divided by zero!

Parth (parthkohli):

911 Emergency, please arrest @SOMU.kvp for dividing by zero!

OpenStudy (anonymous):

hey y not?but logic okay

OpenStudy (anonymous):

how can you divide nothing by nothing

OpenStudy (anonymous):

hey parth i have fear

OpenStudy (anonymous):

Oh yeah The Open Study Area Police Force.....

OpenStudy (anonymous):

I called the police they will arrive in this page soon......

OpenStudy (anonymous):

Congratulations for Green @mukushla

OpenStudy (anonymous):

y not posible? 0 divide 0 nothing friend nothing mean 0

OpenStudy (anonymous):

And I must say looking great in the pic too.. @mukushla

OpenStudy (anonymous):

collisto is the police and he does not approve

OpenStudy (anonymous):

okay okay cool friend just for fun in maths

OpenStudy (anonymous):

For fun you joke with the "Maths"......... LolZZzz.

OpenStudy (anonymous):

Congratulations for Green @mukushla

OpenStudy (anonymous):

And I must say looking great in the pic too @mukushla Ha ha ha ha....

OpenStudy (lgbasallote):

0 = 2 = tw0 = tw(0) = 0 Q.E.D.

OpenStudy (anonymous):

Wow @lgbasallote you are so intelligent..

Parth (parthkohli):

Assume that 0 = 2. Done.

OpenStudy (lgbasallote):

it's called thinking fourth-dimensionally

OpenStudy (anonymous):

There is one more way to prove 0 = 2

Parth (parthkohli):

Assume 0 = 2. Multiply both sides by 0.

OpenStudy (anonymous):

yaa lebasallote

Parth (parthkohli):

0 = 2 0 * 0 = 2 * 0 0 = 0 Proved!

OpenStudy (anonymous):

hey parth what a cute answer!

OpenStudy (anonymous):

LHS : 0 Multiply it with 0: \(0 \times 0 = 0\) -------1 RHS : 2 Multiply it with 0: \(0 \times 2 = 0\) ----------2 From 1 and 2, LHS = RHS And this is called thinking Fifth dimensionally...

OpenStudy (anonymous):

tsk tsk thats only 4th

OpenStudy (anonymous):

think deeper @waterineyes

OpenStudy (anonymous):

For that I have to dig Earth..

OpenStudy (anonymous):

hey nick r u student?

OpenStudy (anonymous):

if im learning im a student

OpenStudy (anonymous):

mmmm i'M A TEACHER

OpenStudy (anonymous):

ninja edit parth lol

OpenStudy (anonymous):

Okay here's a solution: Lets assume we are in the era where there is no mathematician born and I am the first person to be born...... Okay Now I declare......